Bond Company budgets the following purchases of direct materials for the first quarter of the year:

January February March
Budgeted purchases $ 150,000 $ 120,000 $ 90,000

All purchases of direct materials are made on credit. On average, the company pays 80% of its purchases in the month of sales and the remainder in the following month.

Required:

1. For the months of February and March, what are the budgeted cash payments for purchases of direct materials under the assumption that there is no (cash) discount for early payment?

2. For the months of February and March, what are the budgeted cash payments for purchases of direct materials under the assumption that the purchase terms are 2/15, net 30? The company’s policy is to take advantage of all cash discounts for early payment.

3a. Using the purchase terms in Requirement 2, calculate the opportunity cost if Bond does not decide to take advantage of the early payment discount.

3b. Can it be considered good economic policy to take advantage of early payment discounts?
Answer #1

1.

February March
Budgeted Cash Payments $126,000 $96,000
Working:
80% in the same month 96000 72000
Remainder 20% in next month
January Purchase 30000
February Purchase 24000

2.  

February March
Budgeted Cash Payments $     124,080.00 $         94,560.00
Working:
80% in the same month 94080 70560
=120000*80%*98% =90000*80%*98%
Remainder 20% in next month
January Purchase 30000
February Purchase 24000

3.  

Opportunity Cost $          3,360.00
(126000+96000-124080-94560)
Opportunity Cost % 1.54%
(3360/218640)*100

4. Yes